Project

General

Profile

Kaon LT Meetings » History » Revision 708

Revision 707 (Garth Huber, 02/16/2024 05:35 PM) → Revision 708/812 (Alicia Postuma, 02/16/2024 06:10 PM)

{{>TOC}} 

 h1. *KaonLT/PionLT Weekly Meetings* 

 KaonLT/PionLT collaboration analysis meetings. Please add your talk to the meeting ahead of time. Add the newest meeting to the top of the list! 

 --- 

 h2. 2024 Meetings 

 --- 

 h3. February 15, 2024 

 * Garth's Notes attachment:mtg_24feb15.txt 
 * Garth's GitHub meeting Notes attachment:GitHub_mtg_24feb14.txt 
 * Richard 
 ** attachment:KaonLTMeeting2_15_24.pdf attachment:dsig_totstaterror_calc.pdf 
 ** Properly Calculated Statistical Uncertainties 
 *** Ali and I went through and properly implemented statistical uncertainties (double-checked with Garth on Monday). This included updating HMS cerenkov and calorimeter efficiencies that needed to use a "golden singles run" to avoid contamination. A bug in the ratio per bin was fixed so that proper average ratio per bins were found. 
 ** SIMC Aerogel/HGCer geometric cuts were implemented. 
 ** Particle Subtraction Script was updated. Just need pion samples from Ali. 
 ** Beginning Unsep-xsect Systematic Studies 
 ** Continued issues with generating new parameterization 
 *** Unsep xsect fits are starting to look good, but the script to generate improved parameterization generates insane fits. Likely a change to the model is needed. In particular, sigT will need adjustments. 
 * Alicia 
 ** Summary of next steps in asymmetry analysis 
 ** Discussion of missing mass using delta correction: unable to reproduce hcana MMpi - learned that this is expected. Will implement a version of Richard's code from recon_hcana  
 ** Exploring -t dependence of different data and simc files 
 ** Slides: attachment:Asymmetry_Postuma_2024-1.pdf 

 


 --- 

 h3. February 8, 2024 

 * Garth's Notes attachment:mtg_24feb08.txt 
 * Richard 
 ** attachment:KaonLTMeeting2_8_24.pdf 
 ** SIMC Weight 
 *** In order to get Ydata/Ysimc to be even close to unity I needed to divide the SIMC weight by 10^6. I was applying the weight for the histograms but not binned data.  
 ** SIMC MM 
 *** The long radiative tail from simc MM has mostly been resolved by using the correct decay length and properly recalculating the MM using recon_hcana. 
 ** MM PID Analysis 
 *** Played around with MM cuts to see PID of leak through/higher K+ channels. 
 *** Increasing aerogel cut from aero > 3.0 to > 4.5 should resolve some of the issues, but more probably needs to be done. 
 ** MM cut: Yield Comparison 
 *** Using an MM cut of 1.10 < MM < 1.18 produced shockingly consistent yields for Q2=2.115 

 --- 

 h3. February 1, 2024 

 * Garth's Notes attachment:mtg_24feb01.txt 
 * Richard 
 ** attachment:KaonLTMeeting2_1_24.pdf 
 ** Dave’s Delta Adjustments 
 *** Dave's adjustment to HMS delta to correct for the hsxpfp correlation is implimented. 
 ** The weight iteration script is working, but there is an issue with either the SIMC weight or unit charge which is causing me to include a 10^-6 factor in the SIMC weight which messes up the rest of the iteration scripts. 

 --- 

 h3. January 25, 2024 

 * Garth/Richard Notes attachment:mtg_24jan25.txt 
 * Richard 
 ** attachment:KaonLTMeeting1_24_24.pdf 
 ** Issues with dummy runs 
 *** The dtype error for some dummy runs was caused by empty trees when applying the CT cuts. 
 *** The source of these fails cuts was because the CT offset applied at the replay level was skipped for these runs and thus the CT peak was not where the cut expected. 
 ** Weird t-distribution 
 *** Multiple peaks seen in the t-distribution were caused by various particle leakthrough. 
 *** This was cleaned up a bit by implimenting a HMS cer cut > 1.5, but there are still significant particles that need to be cleaned up. 
 ** Fixed binning issue 
 *** Issue with the binning script was causing many average kin per bin values to be zero.  
 *** The scripts were completely rewritten and things look very nice now. 
 ** Separated Xsects Script 
 *** Finished debugging the script that separates the cross section.  
 *** The binning issue was hiding other FORTRAN related issues, in particular, with the input files that are generated. 
 ** SIMC MM 
 *** Very broad and long tail on kaon MM generated by SIMC. Still under investigation. 
 ** SIMC High Q2 
 *** There were negative MM and other variables being generated by SIMC for high Q2 settings. 
 *** Likely caused by sigT going negative. By setting some of sigT's parameterization to zero, the issue should be resolved. Still under investigation. 
 * Dave 
 ** HMS delta-hsxpfp correction for KaonLT data  
 ** attachment:DG_24jan25_hms_hsxpfp.pdf 

 --- 

 h3. January 11, 2024 

 * Garth's Notes attachment:mtg_24jan11.txt 
 * Garth's GitHub meeting Notes    attachment:GitHub_mtg_24jan10.txt 
 * Richard 
 ** attachment:D1_2024WinterHallCTalk.pdf 
 ** Most bad runs (i.e. those replays/analysis files that failed for various regions) for Q2=2.1 have been properly run. There are still a few outstanding runs that need to be further investigated. 
 ** Equations for the xsect iteration script have been properly adjusted to match those found in the fortran script of SIMC (physics_iterate.f). 
 ** Testing with Q2=2.1 is underway with a minor bug with efficiency calculation that needs to be resolved. Once this is done, this setting will test the xsects iteration and separated xsects scripts. 
 * Dave  
 ** HMS delta-hsxpfp correction for P_HMS=5.292GeV/c    attachment:DG_24jan11_hms_hsxpfp.pptx attachment:DG_24jan11_hms_hsxpfp.pdf 

 --- 

 h3. January 4, 2024 

 * Garth's Notes attachment:mtg_24jan04.txt 
 * Nacer: beta SHMS calibration attachment:20240104_kaonLT_nacer.pdf 

 --- 

 h2. 2023 Meetings 

 --- 

 h3. December 21, 2023 

 * Garth's Notes attachment:mtg_23dec21.txt 
 * Richard 
 ** attachment:FinalHeepOffsets_v3.pdf 
 *** Updated heep plots without HGCer cuts, very small change (order of ~0.1 MeV). Also updated with almost final efficiencies (missing Hodo uncern). 
 ** attachment:10p6_Prod_2023_12_18.pdf attachment:8p2_Prod_2023_12_18.pdf attachment:6p2_Prod_2023_12_18.pdf attachment:4p9_Prod_2023_12_18.pdf attachment:3p8_Prod_2023_12_18.pdf 
 *** Plots of efficiencies per energy. 
 ** attachment:heep_yield.pdf 
 *** Heep yield plots without uncertainty. Statistical uncertainty is done, but systematic needs to be finished. 
 * Vijay (updated on Dec 27, 2023 (Heep plot)) 
 ** The final summary of the lumi and elastic studies in summer 2019 data. attachment:PionLT_Smmer19_Lumi_elastic_Studies_Dec21_2023.pdf 
 ** Next Plan: Physics analysis (LT separation) for Q2 = 0.38 GeV2 in the summer 2019 data. 

 --- 

 h3. December 14, 2023 

 * Garth's Notes attachment:mtg_23dec14.txt 
 * Vijay 
 ** The summary of lumi and elastic studies in summer 2019 data. 
 ** Carbon Relative Yield: shows a flat line with the -0.243/100uA% slope. 
 ** LH2 Relative Yield: shows a slope of -5.06/100uA%. 
 ** Average Elastic Cross-Section: determined to 0.9915+-0.0136.  
 ** I've compared the study with previous ones. attachment:PionLT_Smmer19_Lumi_elastic_Studies.pdf 
 ** Next plan: Replay the physics data with determined offsets for the physics analysis.  
 * Nacer: beta SHMS calibration attachment:20231214_kaonLT_nacer.pdf 

 --- 

 h3. December 7, 2023 

 * Garth/Nacer Notes attachment:mtg_23dec07.txt 
 * Richard 
 ** attachment:FinalHeepOffsets_v2.pdf 
 ** No Offsets: "10p6":https://redmine.jlab.org/attachments/2266 "8p2":https://redmine.jlab.org/attachments/2262 "6p2":https://redmine.jlab.org/attachments/2263 "4p9":https://redmine.jlab.org/attachments/2264 "3p8":https://redmine.jlab.org/attachments/2265 
 ** Offsets: "10p6":https://redmine.jlab.org/attachments/2261 "8p2":https://redmine.jlab.org/attachments/2257 "6p2":https://redmine.jlab.org/attachments/2258 "4p9":https://redmine.jlab.org/attachments/2259 "3p8":https://redmine.jlab.org/attachments/2260 
 ** There was an oversight in the BPM params and they were not pushed to the proper git repo. The offsets had to be redone and the new final offsets are implimented. 
 ** Full Replay Pass 3 is underway. 

 --- 

 h3. November 30, 2023 

 * Garth/Nacer Notes attachment:mtg_23nov30.txt 
 * Nathan 
 ** Update on Coin Reference time cuts: attachment:RefTimes_Update_2023-11-30.pdf 
 ** Cuts decided, quality checks on data and implementation to be done 
 * Richard 
 ** attachment:KaonLT_FinalHeepOffsets.pdf 
 ** Finalized KaonLT Offsets (including OOP) 
 ** In the process of including new offsets in a new replay batch. Should be finished by the end of the weekend. 

 --- 

 h3. November 16, 2023 

 * Garth/Nacer Notes attachment:mtg_23nov16.txt 
 * Ali 
 ** attachment:KaonLT_Meeting_Nov_16_2023.pdf 
 ** CoinTime / Beta Leakage Study is now finished. Mark's suggestion to change ptof_tolerance variable from 100 to 2 has fixed the bad beta events. The efficiency for 10.6 and 6.2 / 8.2 GeV Heep dta is almost identical. No need to correct for bad beta / CoinTime events. 
 * Richard 
 ** attachment:KaonLTMeeting11_16_23.pdf 
 ** Fixing the SHMS momentum offset and playing around with some decent offsets for all energies besides 10p6. Final numbers are being determined. 
 * Vijay 
 ** The 'Out of Plane offset' for the summer 2019 data has been comprehensively examined and verified.  
 ** It has proven to be highly effective in addressing the pmy variables of the summer 2019 dataset. attachment:Heep_pmy_summer19_offset.pdf 

 --- 

 h3. November 9, 2003 

 * Garth/Nacer Notes attachment:mtg_23nov09.txt 
 * Richard 
 ** attachment:KaonLTMeeting11_9_23.pdf 
 ** A variety of heep offsets were applied, provided by Garth's global results. It is shown that adapting the 5 beam energy global offsets with varying the SHMS momentum offset allows for the best emiss peak positions. 
 ** OOP offsets are also looked at, but there is an obvious issue with their implementation that needs to be resolved. 
 * Garth 
 ** PMY fit to determine OOP offsets, for 3 different sets of beam energies attachment:offset_pmy2c.pdf 

 --- 

 h3. November 2, 2023 

 * Garth/Nacer Notes attachment:mtg_23nov02.txt 
 * Richard 
 ** attachment:KaonLTMeeting11_2_23.pdf 
 ** Garth's global high energy offsets were applied. There is a shift in emiss with dpp offset of -0.1% 

 --- 

 h3. October 24, 2023 

 * Richard/Tanja/Dave Notes attachment:mtg_23oct24.txt 

 --- 

 h3. October 17, 2023 

 * Garth/Nacer Notes attachment:mtg_23oct17.txt 
 * Richard 
 ** "kaon_xsects_Q2p1W2p95":https://redmine.jlab.org/attachments/2210 
 ** By setting the parameters for the interference terms to zero, the spike in data between 220-300 deg has been removed, and all ratios are within ~10. The weight iteration script is being debugged and thoroughly checked. 
 * Vijay 
 ** The testing of global fitting parameters (offsets) for the summer 2019 data (Garth study) has been completed.  
 ** It has been determined that employing a single set of offsets for all energies within the summer 2019 dataset does not adequately address all variables. 
 ** Nonetheless, the W variable exhibits a strong agreement with the SIMC for energies of 2.7 and 4.5 GeV. However, at 3.6 GeV, a slight over-correction is observed. 
 ** attachment:Vijay_Oct17_23.pdf 

 --- 

 h3. October 16, 2023 

 * "Notes from Cointime Leakage Meeting":https://redmine.jlab.org/attachments/2211 
 ** Attendance: Alicia P., Ali U., Garth H., Richard T. 

 --- 

 h3. October 10, 2023 

 * Garth/Nacer Notes attachment:mtg_23oct10.txt 
 * Richard 
 ** attachment:KaonLTMeeting10_10_23.pdf 
 ** As expected, the bifurcation seen in the CT vs beta (E=10.6) is almost non-existent when kaon PID cuts are applied. This is likely due to the lower statistics compared to what Alicia showed with pion PID. A correction should be applied regardless which will determined by the pion data. 

 --- 

 h3. October 3, 2023 

 * Garth/Nacer/Richard Notes attachment:mtg_23oct03.txt 
 * Richard 
 ** attachment:kaon_xsects_Q2p1W2p95.pdf 
 ** There are some issues with the Right SIMC setting returning zeros, which is part of the cause of the odd values. This is still being debugged, it is unclear if this is a binning or simc issue. 
 * Plots from post-meeting on CoinTime "leakage" 
 ** attachment:cointime_mmpi_10p6.pdf 
 ** attachment:heep_coin_time_10p6.pdf 
 ** attachment:ctime_vs_beta.pdf 
 ** attachment:cointime_beta_10p6.pdf 

 --- 

 h3. September 26, 2023 

 * Garth/Nacer Notes attachment:mtg_23sep26.txt 
 * Richard 
 ** attachment:kaon_unsep_xsects_Q2p1W2p95.pdf 
 ** Fixed issues with the binning. Adapting parameter optimization code for iterating weights of SIMC 
 ** Ratio for low eps of 2.1 looks good (all within 0.1-2.0), but high eps is all over the place. 
 * Alicia 
 ** Explored coin time: interesting feature depends on missing mass 
 ** attachment:log_ctime.pdf dark blue: no MM cut, green: loose MM cut, red: tight MM cut 
 ** attachment:mmcointime.pdf 2D plot of coin time vs MM 

 --- 

 h3. September 19, 2023 

 * Garth/Nacer Notes attachment:mtg_23sep19.txt 
 * Alicia 
 ** Finalizing asymmetry analysis: three unresolved questions on PID/event selection 
 ** attachment:Asymmetry_Postuma_Sep16.pdf 

 --- 

 h3. September 12, 2023 

 * Garth/Nacer Notes attachment:mtg_23sep12.txt 
 * Follow-up meeting on LTsep procedure attachment:LTsep_mtg_23sep13.txt 

 --- 

 h3. September 07, 2023 

 * Garth/Nacer Notes attachment:mtg_23sep07.txt 
 * Vijay 
 ** SHMS hodoscope individual bar efficiency (pion) studied (Efficiency vs Run Number and Efficiency vs Coin trigger Rate) for all the production Runs. attachment:Vijay_Sep07_2023.pdf 
 * Richard 
 ** "HMS Current Carbon":https://redmine.jlab.org/attachments/2118 
 ** "HMS Current LH2":https://redmine.jlab.org/attachments/2119 
 ** Finalized slopes for luminosity analysis from linear regression of HMS electron data from carbon and LH2. Carbon is flat (m0=1.688e-05±4.305e-04) and LH2 has a slope of ~8%+/-2% (m0=-7.899e-04±1.829e-04) 
 * Alicia 
 ** Preliminary Q2 scan: attachment:Scan_prelim.pdf 

 --- 

 h3. August 31, 2023 

 * Garth/Nacer Notes attachment:mtg_23aug31.txt 
 * Richard 
 ** attachment:KaonLTMeeting8_31_23.pdf 
 ** "Regression HMS carbon current":https://redmine.jlab.org/attachments/2072 
 ** "Regression HMS carbon rate":https://redmine.jlab.org/attachments/2074 
 ** "Regression HMS LH2 current":https://redmine.jlab.org/attachments/2071 
 ** "Regression HMS LH2 rate":https://redmine.jlab.org/attachments/2073 
 ** "Regression SHMS carbon current":https://redmine.jlab.org/attachments/2075 
 ** "Regression SHMS carbon rate":https://redmine.jlab.org/attachments/2076 
 ** "Regression SHMS LH2 current":https://redmine.jlab.org/attachments/2077 
 ** "Regression SHMS LH2 rate":https://redmine.jlab.org/attachments/2078 
 ** Luminosity analysis is finally complete. Results are consistent with Carlos Yero! 
 * Alicia 
 ** Binning Monte Carlo: attachment:Asymmetry_Postuma_BinMC2.pdf 
 * Garth 
 ** In-plane offset global constrained fit attachment:results_summary_23aug28.txt 
 ** Out-of-plane offset plot attachment:offset_oop.pdf 

 --- 

 h3. August 24, 2023 

 * Garth/Nacer Notes attachment:mtg_23aug24.txt 
 * Vijay 
 ** All efficiencies for the PionLT summer 2019 data, that are required for the cross-section study, have been studied. attachment:Vijay_update_Aug24_2023.pdf 
 ** Starting soon the elastic cross-section study for the finalization. 


 --- 

 h3. August 17, 2023 

 * Garth's Notes attachment:mtg_23aug17.txt 
 * Alicia 
 ** Several updates: pi-delta BSA and binning for pi-neutron analysis 
 ** attachment:Asymmetry_Postuma_PiDelta.pdf 
 ** attachment:Asymmetry_Postuma_Binning2.pdf 
 ** Also discussion of error analysis: will use unsimplified version of asymmetry error. 
 * Nathan 
 ** Slides from conlsion of hgcer calibration https://redmine.jlab.org/attachments/2063 

 --- 

 h3. August 10, 2023 

 * Garth's Notes attachment:mtg_23aug10.txt 
 * Richard 
 ** attachment:KaonLTMeeting8_10_23.pdf 
 ** t/phi binning of kinematics is looking a bit odd. The procedure is broken down here. 
 * Alicia 
 ** Working on binning data for pi-neutron asymmetry 
 ** attachment:Asymmetry_Postuma_Binning1.pdf 

 --- 

 h3. July 27, 2023 

 * Garth's Notes attachment:mtg_23jul27.txt 

 --- 

 h3. July 20, 2023 

 * Garth's Notes attachment:mtg_23jul20.txt 
 * Vijay 
 ** The HMS calorimeter efficiency study is completed from all the ‘Heep coin data’ (3 beam energy) of the summer 2019 run period. attachment:Vijay_July20_2023.pdf 
 ** The final number of this is *{color:green}@0.9979 ± 4.004e-05@*.   
 * Richard 
 ** "kaon_FullAnalysis_Q5p5W3p02_lowe":https://redmine.jlab.org/attachments/2028 "kaon_FullAnalysis_Q5p5W3p02_highe_no_hgcer_cut":https://redmine.jlab.org/attachments/2030 "kaon_FullAnalysis_Q5p5W3p02_highe":https://redmine.jlab.org/attachments/2029 
 ** Peter's kaon PID suggestions are implemented, and the samples are almost there. These include aero>4, RF>+/-0.5, hgcer hole cut. Note: these are with just the spring hole cut (for testing). The proper fall hole cut will be added soon. 

 --- 

 h3. July 13, 2023 

 * Garth's Notes attachment:mtg_23jul13.txt 

 --- 

 h3. July 6, 2023 

 * Garth's Notes attachment:mtg_23jul06.txt 
 * Vijay 
 ** The HMS Calorimeter efficiency study from all the production data at Q2 = 0.38 GeV. attachment:Vijay_July6_2023.pdf 

 --- 

 h3. June 28, 2023 

 * Garth's Notes attachment:mtg_23jun28.txt 
 * Richard's Notes from meeting with Peter attachment:peter_mtg_23jun25.txt 
 * Vijay 
 ** EDTM Live Time testing for the Lumi study from summer 2019 data is done. 
 ** attachment:Vijay_KaonLT_meeting_Jun28_2023.pdf 
 * Richard 
 ** attachment:KaonLTMeeting6_28_23.pdf 
 ** Incorporated Peter (and other's) bcm correction. Some things look better, but there are issues with some TLT values, which is throwing yields off.  

 --- 

 h3. June 22, 2023 

 * Garth's Notes attachment:mtg_23jun22.txt 
 * Richard 
 ** attachment:KaonLTMeeting6_22_23.pdf 
 ** Adjusted aerogel cut for kaons to 3.0 from 1.5 because of a suggestion by Peter. This eliminates a bit more protons that slip through. 

 --- 

 h3. June 15, 2023 

 * Dave's Notes attachment:mtg_23jun15.txt 

 --- 

 h3. June 8, 2023 

 * Dave's Notes attachment:mtg_23jun08.txt 
 * Richard 
 ** attachment:KaonLTMeeting6_8_23.pdf 
 ** Linear regression for production also includes higher rates (>1kHz) and things still look fine. 
 ** LH2 needs cuts to be adjusted, things are pretty off for now.  
 * Alicia 
 ** attachment:Asymmetry_Postuma_Updates_2023.pdf 
 ** Errors, models for BSA analysis 
 * Nathan 
 ** Finished Noble gas calibration attachment:CalibUpdate_2023-06-15.pdf 
 ** Uploaded slides from previous update, with a few extra plots that I had  

 --- 

 h3. June 1, 2023 

 * Dave's Notes attachment:mtg_23jun01.txt 
 * Richard 
 ** attachment:KaonLTMeeting6_1_23.pdf 
 ** Linear regression for coincidence runs were compared to the singles luminosity runs. The slope for the coincidence is approximately an order of magnitude better than the singles. 

 --- 

 h3. May 25, 2023 

 * Garth's Notes attachment:mtg_23may25.txt 
 * Vijay 
 ** HMS Calorimeter efficiency study at Q2 = 0.38 GeV2, for all production and heep coin data. attachment:Vijay_May25_2023.pdf 
 ** We discussed in our group meeting that it should be ok to take error weighted average of all the heep coin data and then correct the production data.  

 * Alicia 
 ** Missing mass by bins - added delta: attachment:mm_simcdata2.pdf 

 --- 

 h3. May 18, 2023 

 * Garth's Notes attachment:mtg_23may18.txt 
 * Ali 
 ** Out of Plane offset study from KaonLT Heep Coindata using Tanja's method. 
 ** attachment:KaonLT_Meeting_May_18.pdf 
 * Vijay 
 ** Raster checks of the summer 2019 data. attachment:Raster_check_summer19.pdf 
 ** I checked the Raster for the summer 2019 data. It looks ok to me. I did not see any slightly tilting issues that Ali saw earlier. 
 ** I do not think that we will need to make any changes to the Raster in the summer 2019 data.  
 ** Please check the attached file and let me know if you have any concerns of the study.  
 * Richard 
 ** attachment:hms_regression_current_carbon.pdf ,attachment:hms_regression_rate_carbon.pdf , attachment:shms_regression_current_carbon.pdf , attachment:shms_regression_rate_carbon.pdf 
 ** I did a least square linear regression on all the luminosity data. There is a good consistency between all data settings and even HMS and SHMS. 

 --- 

 h3. May 11, 2023 

 * Dave's Notes attachment:mtg_23may11.txt 
 * Richard 
 ** attachment:KaonLTMeeting5_11_23.pdf 
 ** Carlos and I found the error in the HMS scaler rate and slope. I was previously doing some testing of ELREAL vs ELCLEAN and I never reverted my testing from ELCLEAN. Once I changed the variable back to ELREAL, the rates and trends match. 
 * Vijay 
 ** HMS Calorimeter efficiency study. attachment:Vijay_May11_2023.pdf 
 ** It looks like the efficiency is less from the production data (Q2 = 0.38 GeV2, High e). 
 ** I'll take a look at the heep coin data in this run period of the experiment and get the efficiency.  
 ** Hopefully that's number, the efficiency from heep coin data, will be reliable to correct the inefficiency of the detector and correct the data run by run. 

 * Ali 
 ** Raster Y sign test. Original sign was working fine. 
 ** attachment:KaonLT_Meeting_May_11.pdf 

 * Alicia 
 ** Updates on BSA analysis 
 ** attachment:Asymmetry_Postuma_May1.pdf 

 --- 

 h3. May 4, 2023 

 * Garth's Notes attachment:mtg_23may04.txt 
 * Richard 
 ** attachment:KaonLTMeeting5_4_23.pdf 
 ** Most of the bad runs have been resolved although there are still a few issues I am sorting out. I also have some plots of the boiling corrections that need to be updated. 
 ** I talked with Jacob last week and he checked my luminosity calculations and they seem fine. Carlos and I are meeting tomorrow at 11 am to try and figure out our discrepancies. 
 ** Found a possible cut issue with the proton peak where cuts were not being properly applied. This is still early in checks. 

 * Ali 
 ** BPM calibration check. Using new file for Spring 2019 data. 
 ** attachment:KaonLT_Meeting_May_4.pdf 

 * Alicia 
 ** Missing mass by bins - added SIMC: attachment:mm_simcdata1.pdf 

 --- 

 h3. April 27, 2023 

 * Garth/Stephen Notes attachment:mtg_23apr27.txt 
 * Vijay 
 ** Updates of the analysis are attached. attachment:Vijay_Apr27_2023.pdf 
 * Richard 
 ** attachment:KaonLTMeeting4_27_23.pdf 
 ** There are still issues with the TLT for some carbon lumi runs. Jacob and I are meeting tomorrow to discuss things. 
 ** Pion subtraction have improved things a lot, but there seems to be issues with the proton subtraction. Adjusted cuts are rerunning and will hopefully help. 
 
 --- 

 h3. April 20, 2023 

 * Garth/Stephen Notes attachment:mtg_23apr20.txt 
 * Nathan - Presented updates on the aerogel calibration. attachment:AerogelCalib_2023-04-20.pdf 
 ** looks like the calibration has significantly improved the separation between signal and zeros. 
 ** for completeness here are all the presentations that Petr gave on the aerogel calibration. attachment:presentation-2020-09-23-kaonlt.pdf attachment:presentation-2020-09-23-kaonlt-summary.pdf 

 --- 

 h3. April 13, 2023 

 * Garth's Notes attachment:mtg_23apr13.txt 
 * Vijay 
 ** The details of the update are attached here. attachment:Vijay_Apr13_2023.pdf 
 * Ali 
 ** Pion PID study for high Q2 KaonLT data. 
 ** Spring 2019 data (low e) attachment:KaonLT_Meeting_Apr_13.pdf 

 --- 

 h3. April 6, 2023 

 * Garth/Stephen Notes attachment:mtg_23apr06.txt 
 * Richard 
 ** attachment:KaonLTMeeting4_6_23.pdf 
 ** "10p6 c #1":https://redmine.jlab.org/attachments/1841 "10p6 LH2 #1":https://redmine.jlab.org/attachments/1843 
 ** "10p6 c #2":https://redmine.jlab.org/attachments/1840 "10p6 LH2 #2":https://redmine.jlab.org/attachments/1844 
 ** "10p6 c #3":https://redmine.jlab.org/attachments/1842 "10p6 LH2 #3":https://redmine.jlab.org/attachments/1845 
 ** "8p2 c #1":https://redmine.jlab.org/attachments/1838 "8p2 LH2 #1":https://redmine.jlab.org/attachments/1839 
 ** Luminosity plots are looking pretty good for LH2 and Carbon. There are a few runs with really bad TLT that need to be resolved. 
 ** I still need to meet with Carlos to figure out HMS scaler issue. 
 ** In the process of subtracting proton events from MM for kaon to try and clean up the right setting (update next week). 

 --- 

 h3. March 30, 2023 

 * Garth/Stephen Notes attachment:mtg_23mar30.txt 
 * Vijay 
 ** Switched the pion routine in SIMC for the pion cross-section. 
 ** Re-ran SIMC to simulate the p(e, e’π+)n reaction for all the settings of Q2 = 0.38 GeV2 data. 
 ** More details are in the talk. attachment:Vijay_KaonLT_meeting_Mar30_2023.pdf 

 --- 

 h3. March 23, 2023 

 * Garth's Notes attachment:mtg_23mar23.txt 
 * Carlos 
 ** relative yields rate dependence 
 https://hallcweb.jlab.org/DocDB/0012/001209/001/rate_dependence_slides_cyero_Mar09_2023.pdf 
 ** event types definitions 
 https://hallcweb.jlab.org/DocDB/0012/001208/001/event_type_definitions_COIN_DAQ_cyero_Mar16_2023.pdf 
 * Ali 
 ** Pion PID study for high Q2 data. This update shows Q2 = 2.1, low e (center) setting. attachment:KaonLT_Meeting_Mar_23.pdf 
 * Richard 
 ** attachment: 
 ** Carlos and I discussed some oddities in a comparison of our lumi scaler results for the HMS. 
 ** "Added quick script to check cuts applied in ltsep python package.":https://github.com/JeffersonLab/UTIL_KAONLT/blob/spring_2022/check_cuts.sh This requires the most update version of the ltsep package, but can be easily downgraded. (just let me know if interested) 
 ** Alicia, Ali, and I are meeting tomorrow at 2:30 to discuss the odd right setting peaks. 

 --- 

 h3. March 16, 2023 

 * Garth/Stephen Notes attachment:mtg_23mar16.txt 
 * Richard 
 ** attachment:KaonLT_Meeting_3_16_23.pdf 
 ** Carlos has been working on the rate dependence issue for the luminosity scans that Casey and I have been seeing. He came to a similar conclusion as me that there will need to be a correction factor that we all agree on that will need to be applied. We are meeting on Monday to discuss things further. "See his slides on the docDB":https://hallcweb.jlab.org/doc-private/ShowDocument?docid=1206 
 ** I think I have tracked down the odd issues with right setting that I have been seeing for most energies. I am not sure why, but it seems that something is going wrong with the replays. I tried rerunning them, but the issue has persisted. My initial guess is that it may be an issue with the TDC cuts, since the coin time distributions were looking off previously.  
 * Nathan 
 ** presented recent updates on the Aerogel calibration: attachment:AerogelCalib_2023-03-16.pdf 
 ** will make multiple calibrations based on PMT4+ dependance 

 --- 

 h3. March 09, 2023 

 * Garth/Stephen Notes attachment:mtg_23mar09.txt 
 * Nathan 
 ** Presented the an update to the aerogel calibration, including a 3 year old presentation from Petr 
 *** presentation attachment:presentation-2020-09-23-kaonlt-summary.pdf 
 ** Also working on the anomalous LD- setting. Things not making sense. Requesting back up. 

 --- 

 h3. March 02, 2023 

 * Garth's Notes attachment:mtg_23mar02.txt 
 * Richard 
 ** attachment:KaonLTMeeting3_2_23.pdf 
 ** Heep SIMC to hcana script is working for missing momentum and energy variables. "Heep 10p6":https://redmine.jlab.org/attachments/1763 "Heep 8p2":https://redmine.jlab.org/attachments/1759 "Heep 6p2":https://redmine.jlab.org/attachments/1760 "Heep 4p9":https://redmine.jlab.org/attachments/1762 "Heep 3p8":https://redmine.jlab.org/attachments/1761 
 ** Fortran scripts for finding average kinematic values and cross sections is working. "Q2=5.5, W=3.02, high eps":https://redmine.jlab.org/attachments/1767 "Q2=5.5, W=3.02, low eps":https://redmine.jlab.org/attachments/1766 
 ** My yield ratios are looking a very weird, more investigation is required. "Ratio and xsects for Q2=5.5, W=3.02, low eps":https://redmine.jlab.org/attachments/1765 
 * Vijay 
 ** MMpi cut updated. 
 ** 8th t-bin adjusted. 
 ** Events per t-bin calculated for each epsilon setting. 
 ** Experimental Yield calculation in each bin is done. 
 ** More details of the update are in the slides. attachment:KaonLT_meeting_March02_2023.pdf 

 --- 

 h3. Feb 23, 2023 

 * Garth/Stephen Notes    attachment:mtg_23feb23.txt 
 * Ali 
 ** Update on HMS Cherenkov and Calorimeter efficiency study. attachment:KaonLT_Meeting_Feb_23.pdf 
 ** Looking at wide range of Heep Singles settings from KaonLT data (primarily low e data). 

 --- 

 h3. Feb 16, 2023 

 * Garth/Stephen Notes attachment:mtg_23feb16.txt 
 * Vijay 
 ** Updated the diamond cut for Q2 = 0.38 GeV2 data. 
 ** t resolution checked from the center setting of SIMC at low epsilon. 
 ** Binned the data in t (total 8 bins). 
 ** t resolution checked for each t bin too. 
 ** Working to bin the data in Phi (16 bins). 
 ** Working to make the Fortran scripts (average_ratios.f, average_kinematics.f & calc_xsect.f) for the summer 2019 data. attachment:KaonLT_meeting_Feb16_2023.pdf 


 --- 

 h3. Feb 9, 2023 

 * Garth/Stephen Notes attachment:mtg_23feb09.txt 
 * Ali 
 ** attachment:KaonLT_Meeting_Feb_09.pdf 
 ** HMS Cerenkov and Calorimeter Study for different kinematics settings. 
 * Richard 
 ** attachment:KaonLTMeeting2_9_23.pdf 
 ** Fixed the odd right distribution for Q2=5.5 by applying more strict cointime cuts. 
 ** SIMC c++ code is finishing up. It currently just updates pmx, pmy, pmz, pm, and em with the hcana version of calculations. Angles will be the next step once things are finished up. 

 --- 

 h3. Feb 2, 2023 

 * Garth/Stephen Notes attachment:mtg_23feb02.txt 
 * Richard 
 ** attachment:KaonLTMeeting2_2_23.pdf 
 ** The simc portion of the LT_analysis code is working but there is a bug I am sorting out which produces incorrect plots 
 ** During a discussion Ali, Vijay, and I had, it was realized that another script is likely missing which I will discuss with Bill. This script is one that combines the phi settings for the fortran code. 
 ** Dave and I met to go over the fixed variable implementation in SIMC. We fixed a few issues which I am in the process of fixing. 
 * Vijay 
 ** Update from the analyzed low epsilon data of Q2 = 0.38 GeV2. attachment:KaonLT_meeting_Feb2_2023.pdf 
 * Alicia 
 ** Preliminary asymmetry results: Q2=3/W=2.32, Q2=4.4/W=2.74, Q2=2.115/W=2.95, aluminum 
 ** attachment:Asymmetry_Postuma_Feb2.pdf 


 --- 

 h3. Jan 26, 2023 

 * Stephen's Notes attachment:mtg_23jan26.txt 
 * Richard 
 ** attachment:KaonLTMeeting1_26_23.pdf attachment:FullAnalysis_Q5p5W3p02_highe.pdf 
 ** LT_analysis script is finishing up. I have the SIMC portion just about working now. I have some other things I am adding because of my conversation with Bill last week. 
 ** SIMC vs Data plots for Q2=5p5 show a very small SIMC. Need to investigate further. 
 * Nathan 
 ** attachment:RefTimes_Update_2023-01-26.pdf  
 ** Asked Carlos Yero about the cointime ref times, he sent me an older email chain attachment:ref_times_email_exchange.txt 
 *** only SHMS variables are used. 
 ** Presented ref times plots, and suggested cuts. 
 *** Cuts are already set up and pull request is sent to LTSep_Analysis_2022 branch 

 --- 

 h3. Jan 19, 2023 

 * Garth/Stephen Notes    attachment:mtg_23jan19.txt 
 * Ali 
 ** Looking at HMS Calorimeter and Cherenkov Efficiency and PID. attachment:KaonLT_Meeting_Jan_19.pdf 
 * Vijay 
 ** attachment:Vijay_Jan19_2023.pdf 
 ** Trying to understand the pion cross-section across the acceptance.  
 * Alicia 
 ** attachment:Asymmetry_Postuma_Updates-6.pdf 
 ** Added dummy target subtraction to asymmetry calculation 
 ** Discussed fitting to extract cross-section ratio 
 ** Requested: details on error calculations: attachment:Asymmetry_Errors.pdf 
 * Richard 
 ** attachment:KaonLTMeeting1_19_23.pdf 
 ** Implemented Jacob's diamond cut script into the setup before Bill's code. This also does the t-binning as well. 
 ** There was missing code for the SIMC implementation which has been found. I am meeting with Bill tomorrow 1/20 to discuss this implementation. Once this is complete, everything is in place to run the Fortran code to get average kinematics and unsep cross sections. 
 * Nathan 
 ** Started working on reference times 
 ** has some extra peaks in some runs 
 *** Should be explained by the triggers on in each run 
 ** attachment:RefTimes_Update_2023-01-19.pdf 

 ---- 

 h3. Jan 4, 2023 

 * Garth's Notes attachment:mtg_23jan04.txt 
 * Ali 
 ** attachment:Jan4_KLT_Meeting.pdf 
 ** All Kaon_LT data has been replayed over the holidays. Slide show list of major changes for the replay. 

 --- 

 h2. 2022 Meetings 

 --- 

 h3. Dec 21, 2022 

 * Garth/Stephen Notes attachment:mtg_22dec21.txt 
 * Richard 
 ** attachment:KaonLTMeeting12_21_22.pdf 
 ** SIMC code was updated to match what is in hcana, using hcana and parts of Carlos' code as a template. Things do not look right though.  
 ** We are a month behind at this point and Bill has finally responded to my questions so I am going to fully concentrate on getting cross sections before the Hall C collaboration meeting in January so that Nathan can present them. I am setting a soft deadline of January 6th to get him early results but will continue to push for cross sections by the talk. 

 --- 

 h3. Dec 14, 2022 

 * Garth/Stephen Notes attachment:mtg_22dec14.txt 

 --- 

 h3. Dec 7, 2022 

 * Garth/Stephen Notes attachment:mtg_22dec07.txt 
 * Ali 
 ** New error calculation (binomial) for efficiencies and Live Times. attachment:KaonLT_Dec_5.pdf 
 * Richard 
 ** "Q2=3.0, W=3.14, lowe":https://redmine.jlab.org/attachments/1707 "Q2=5.5, W=3.02, lowe":https://redmine.jlab.org/attachments/1708 "Q2=4.4, W=2.74, lowe":https://redmine.jlab.org/attachments/1709 "Q2=4.4, W=2.74, highe":https://redmine.jlab.org/attachments/1710 "Q2=5.5, W=3.02, highe":https://redmine.jlab.org/attachments/1711 "Q2=3.0, W=3.14, highe":https://redmine.jlab.org/attachments/1712 
 ** A new subroutine has been added for the hcana variable calculations in SIMC, this is ongoing. 
 ** Plots of right/left/center for each physics setting have been added, including efficiencies and t-bin plots (no diamond cuts yet!). These plots show the input to Bill's code for comparison. These need to be updated with Ali's updates so a new set of replays will need to be analyzed. 
 ** Bill is doing a write-up with some questions I had which will hopefully fix the remaining issues I've been having with his code. 
 * Alicia: attachment:Asymmetry_Postuma_Dec07.pdf 

 --- 

 h3. Nov 30, 2022 

 * Garth/Stephen Notes attachment:mtg_22nov30.txt 

 --- 

 h3. Nov 23, 2022 

 * Garth/Stephen Notes attachment:mtg_22nov23.txt 
 * Alicia asymmetry update attachment:Asymmetry_Postuma_Updates.pdf 

 --- 

 h3. Nov 16, 2022 

 * Garth/Stephen Notes attachment:mtg_22nov16.txt 
 * Richard 
 ** "8p2 LH2":https://redmine.jlab.org/attachments/1684 "8p2 C":https://redmine.jlab.org/attachments/1685 
 ** "10p6 C 2":https://redmine.jlab.org/attachments/1686 "10p6 C 1":https://redmine.jlab.org/attachments/1687 "10p6 LH2 1":https://redmine.jlab.org/attachments/1688 "10p6 C 3":https://redmine.jlab.org/attachments/1689 "10p6 LH2 2":https://redmine.jlab.org/attachments/1690 "10p6 LH2 3":https://redmine.jlab.org/attachments/1691 
 ** Fits for lumi scans look great for HMS across the board. SHMS no track also look pretty good but tracking still needs work. Ali has adjusted the tracking parameters in the replay which should help things. Signs of boiling are also now evident in HMS for some runs. 
 ** Odd issue with the yields going on that I need to discuss with Bill. It seems to be a cut discrepancy that is eliminating events, I may be missing a cut that he has hidden somewhere. Besides this, the cross-section code is being edited for HeeP so we can get some cross-section plots on this soon. 
 ** I found a script for raster calibrations. The script is being fixed as it is currently a bit of a mess. 
 ** https://redmine.jlab.org/projects/kltexp/wiki/Hall_C_Raster_Algorithm 
 * Stephen 
 ** Updates to the hcana RFTime class 
 ** Added in particle dependent corrections for e/pi/K/p for each spectrometer 
 *** Calculate beta for each particle type (from momentum, delta) 
 *** Determine path length from xptar and delta 
 *** Calculate ToF 
 *** Add this in modulo that determines RFtime 
 *** Tried method Shuo used initially, didn't really do anything 
 ** Some further updates needed before it will be pushed to the hallc repo 
 ** "Slides":https://redmine.jlab.org/attachments/1694  
 ** "RF corrections applied exactly as Shuo applied them":https://redmine.jlab.org/attachments/1696 "RF corrections determined as Shuo did, but applied inside the modulo fn":https://redmine.jlab.org/attachments/1697  
 ** "RF correction simply the ToF added in the modulo fn, seems to work":https://redmine.jlab.org/attachments/1695 
 * Vijay (presented by GH) 
 ** low Q2 attachment:CPULT_lowQ2.png attachment:TLT_lowQ2.png attachment:ELT_lowQ2.png 

 --- 

 h3. Nov 09, 2022 

 * Garth/Stephen Notes attachment:mtg_22nov09.txt 
 * Bill (Wenliang) Li L/T-separation code tutorial. 
 ** Video recording is posted at: http://lichen.phys.uregina.ca/~huberg/kaon/LT_tutorial_20221109_1920x1080.mp4 
 * Vijay 
 ** The scaler Yield looks very nice at 2.7 GeV PionLT data. Below 10 uA could be fixed by placing some of the offsets in the BCM calibration. attachment:Vijay_LumiStuy.pdf  

 --- 

 h3. Nov 2, 2022 

 * Garth/Stephen notes attachment:mtg_22nov02.txt 
 * Vijay 
 ** The first round of the Lumi study in 2.7 GeV data of the PionLT experiment was preceded. The plots are attached, but I do not have any conclusions for now. Please see the attachment. 
 attachment:Vijay_REYIELD_2p7C.pdf  
  
 * Ali 
 ** Heep Coin Offset study using Peter Bosted's correction factor for HMS central momenta. The results from 8.2 and 10.6 GeV improved after applying the offset whereas 6.2 over corrects the distributions as they are already looking good with no offset. Data and SIMC comparison look very good for W distribution but doesn't perfectly match for missing variables. 
 attachment:KaonLT_Meeting_Nov_2.pdf 
 * Jacob 
 ** Regarding Ali's work on the Pc offsets at high momentum, I referenced the delta corrections I made early last year. In EPICS, the NMR did not appear to be locked and in the middle of the scan, the NMR was changed. Slide 15 has the offsets to the central momentum, Pc. 
 attachment:HMS_Optimization_at_High_Momentum.pdf 
 * Richard 
 ** "6p2 C":https://redmine.jlab.org/attachments/1658 "6p2 LH2":https://redmine.jlab.org/attachments/1659 
 ** "10p6 C 1":https://redmine.jlab.org/attachments/1662 "10p6 C 2":https://redmine.jlab.org/attachments/1663 "10p6 C 3":https://redmine.jlab.org/attachments/1664 "10p6 LH2 1":https://redmine.jlab.org/attachments/1665 "10p6 LH2 2":https://redmine.jlab.org/attachments/1666 "10p6 LH2 3":https://redmine.jlab.org/attachments/1667 
 ** Luminosities are looking very nice. These now include the ones Peter pointed out that were taken during SIDIS. Nearly all e-e runs are under 5%. 
 ** I have debugged all of the issues I was having with Bill's code. I am now in the process of merging the two frameworks which is mainly just making names, conventions, etc. consistent. 
 ** I have been writing up some questions for Bill next week when he presents. I'd like to also have these two frameworks pretty much merged by then. 

 
 ---- 

 h3. Oct 26, 2022 

 * DaveG notes attachment:mtg_22oct26.txt 

 ---- 

 h3. Oct 19, 2022 

 * Garth/Stephen notes attachment:mtg_22oct19.txt 
 * Richard 
 ** The trend in the SHMS track yields is still present in the positive polarity luminosity runs. 
 ** Talk for the Hall C Quarterly meeting was discussed.  

 ---- 

 h3. Oct 12, 2022 

 * Garth/Stephen notes attachment:mtg_22oct12.txt 

 ---- 

 h3. Oct 4, 2022 

 * Garth/Stephen notes attachment:mtg_22oct05.txt 
 * Richard 
 ** attachment:KaonLTMeeting10_4_22.pdf 
 ** Fixed some issues with the luminosity scripts with updated replay and analysis script. The biggest issue was the really low beam on times which came from BCM4a being called by accident in some cases instead of BCM1. 
 *** It should be noted, there are huge discrepancies between BCM4A and BCM1 in some cases (like 20+ uA). 
 ** Ran into some issues with the cross-section code. I am in debug mode trying to find the cause. 
 ** Tanja sent me some awesome fpi2 raster correction documents as well as procedures for verifying the beam and target positions. I am in the process of reading through/beginning these procedures. A Redmine wiki page will be linked when completed. 

 ---- 

 h3. Sept 28, 2022 

 * Garth/Stephen notes attachment:mtg_22sep28.txt 
 * Jacob 
 ** HMS optics fitting - attachment:HMS_Optics_Calibrations_Update.pdf 

 ---- 

 h3. Sept 21, 2022 

 * Garth/Stephen notes attachment:mtg_22sep21.txt 
 * Richard 
 ** attachment:Yield_6p2cl1.pdf attachment:Yield_10p6cl2.pdf attachment:Yield_10p6cl1.pdf 
 ** Lumi SHMS tracking is still showing the trend from 9/8. It was found that the lumi replay script is not updated with many of Ali and Richard's most recent changes. These changes are being applied and a set of new replays will be done. 
 ** Bill's cross-section code is broken into two major parts. The root analysis part, which takes all the yields, uncertainties, etc., is nearly complete with the connecting ltsep framework script finishing up. The second part, where many of the final plots and cross-sections are done, is a hard-coded mess. Richard is in the process of making this more dynamic. This aspect will take more time but the root analysis part is the bulk of the analysis anyways. 
 * Ali 
 ** attachment:Ali_Sep21_Analysis.pdf 
 ** Heep Coin analysis for high Q2 KaonLT data (6.2, 8.2, 10.6 GeV beam energy) has now all the efficiencies/live times applied. The results for all 3 beam energies are looking very good. Data/SIMC ratios are agreeing very well (< 3%) and the shapes are matching up nicely. Next step is to finalize offsets for both low and high Q2 KaonLT data. Also proton absorption correction and target boiling needs to be applied. 

 ---- 

 h3. Sept 8, 2022 

 * Garth/Stephen Notes attachment:mtg_22sep08.txt 
 * Ali 
 ** attachment:Sep8_Analysis.pdf 
 ** Heep Coin Analysis update. Fixed the issue with emiss and pmiss. Includes number of debug tests. 
 ** attachment:FullAnalysis_4p9-2.pdf attachment:FullAnalysis_8p2.pdf attachment:FullAnalysis_6p2.pdf attachment:FullAnalysis_3p8-1.pdf attachment:FullAnalysis_10p6-1.pdf 
 ** Full analysis files are the latest data/simc comparison for different beam energies (No offsets applied yet)  
 * Richard 
 ** attachment:KaonLTMeeting9_8_22.pdf 
 ** attachment:Yield_6p2_l1_all.pdf attachment:Yield_10p6_l1_all.pdf attachment:Yield_10p6_l2_all.pdf 
 ** Improved total uncertainties 
 ** In-depth with tracking cut analysis underway 
 ** Poor TLT consistent with poor CPULTs 

 ---- 

 h3. Aug 31, 2022 

 * Garth/Stephen Notes attachment:mtg_22aug31.txt 
 * Richard 
 ** attachment:KaonLTMeeting8_31_22.pdf 
 ** Improvements to luminosity analysis. Things are getting within a few percent. There are still some obvious issues with dependency leaking into the cuts, especially for tracking which is being investigated further. 
 ** Uncertainties for the luminosities and heep are next up. 

 ---- 

 h3. Aug 26, 2022 

 * Meeting on L/T-separation procedure 
 ** Garth's slides attachment:LT_sep_iterations.pdf 
 ** Bill's tutorial (Nov 2018) attachment:Kaon_LT_tutorial_18nov28.pdf 
 ** Stephen's notes attachment:LTsep_mtg_22aug26.txt 
 ** Recording of Sept 27 repeat at: http://lichen.phys.uregina.ca/~huberg/kaon/LT_tutorial_20220927_1920x1080.mp4 

 ---- 

 h3. Aug 17, 2022 

 * Garth/Stephen Notes attachment:mtg_22aug17.txt 
 * Vijay 
 ** The summary of the Heepcheck study is attached here. attachment:Vijay_Aug17_2022.pdf  
 ** 3.6 and 3.9 GeV data do NOT have the angle offsets. This has been established in the study. 
 ** I've plotted the offsets vs SHMS central momentum. Plots are attached in the attached pdf file. 
 * Richard 
 ** The 10.6 GeV heep analysis is going well but the broad momentum distributions are making offsets difficult at times. 
 ** I have decided to give a second pair of eyes to Ali's hcana vs simc calculation investigation. It is concluded that the broad peaks of the hcana are likely caused by the secondary particle calculation which is gathered from the spectrometer track info.  
 ** Corrections to the simc pmiss components have been added to the fortran code which corrects for a minus sign that is in hcana but missing from simc. 
 ** attachment:FullAnalysis_Offset_10p6.pdf 
 ** attachment:HcanavSimc_scatter_calc.pdf 
 

 --- 

 h3. Aug 4, 2022 

 * Garth/Stephen Notes attachment:mtg_22aug04.txt 
 * Vijay 
 ** The offset HeepCheck study at 3.9, 4.9 (KaonLT), 2.7, 3.6 and 4.5 (PionLT) GeV has been completed. attachment:Vijay_HeepCheckStudy.pdf   
 * Richard 
 ** attachment:KaonLTMeeting_8_4_22.pdf 
 ** HeeP coin for 10p6 setting is looking pretty good. I still need to make some minor adjustments. 
 ** TLT for Lumi is looking much better. The yields also are looking better but cut adjustments need to be made. 
 ** HeeP singles have some odd issues that need to be resolved before offset analysis can begin. 
 ** Hodoscope calibrations are all done except 6p2. 

 --- 

 h3. July 14, 2022 

 * Garth/Stephen Notes attachment:mtg_22jul14.txt 
 * Richard 
 ** attachment:KaonLTMeeting7_14_22.pdf 
 ** The W peaks for SIMC and data are very far offset from one another. This is being investigated. 
 * Vijay 
 ** Heep Data MC comparo w/o offsets attachment:Heep_coin_no_offsets_vijay_22jul08.pdf 

 ---- 

 h3. June 28, 2022 

 * Garth/Stephen Notes attachment:mtg_22jun28.txt 
 * Richard 
 ** Aerogel XY plots vs NPE to check out the tray location and efficiencies. 
 ** Hodoscope and calorimeter scripts working. Hodoscope calibration looks good for 10p6 but needs improvement for other beam energies. 
 ** HGCer efficiency and plots added to efficiency script. Need to discuss with Vijay on specifics of cuts and efficiencies. 
 ** Started looking in detail at Bill's code. Some scripts and links are missing/broken. 
 * Jacob 
 ** EDTM Study Report completed, with newer versions of the TLT equations attachment:EDTM_Study_Report.pdf 
 ** Graphical Guide for Event Types created to accompany the report, illustrating prescaling with EDTM attachment:Graphical_Guide_Evt_Type.pdf 
 ** Plan to update report file template with new equations and errors soon 

 ---- 

 h3. June 14, 2022 

 * Garth's Notes attachment:efficiencies_22jun13.txt 
 * Richard 
 ** Fixed issues with simc script and 10p6 Heepcoin data. Issue with data was an offset of 40ns that was corrected halfway through runs. 
 ** Aerogel XY plots vs NPE to check out the tray location and efficiencies. 
 ** Efficiency plots for all production runs. 
 ** attachment:KaonLTMeeting6_14_22.pdf 

 ---- 

 h3. May 31, 2022 

 * Garth and Stephen's Notes attachment:mtg_22may31.txt attachment:efficiencies_22jun02.txt 
 * Vijay  
 ** The updated Heep coin study (Heepcheck program) at 3.9 GeV KaonLT data is attached here attachment:Vijay_KaonLT_meeting_May31_2022.pdf . 
 * Ali 
 ** First look at the HEEP COIN Data/SIMC comparison for the 8.2 GeV Kaon-LT data attachment:KaonLT_May31.pdf 

 ---- 

 h3. May 17, 2022 


 * Garth and Stephen's Notes attachment:mtg_22may17.txt attachment:efficiencies_22may25.txt 
 * Richard 
 ** Updated heep coin scripts for the farm. Ali is testing them and should be able to get plots soon. 
 ** There was a small typo in the simc input file that dramatically changed the shape of the plots. attachment:KaonLTMeeting5_17_22.pdf 
 ** Updated version of efficiency columns for the script that will grab efficiencies and various other values. attachment:efficiency_output_list.docx 
 * Vijay  
 ** The updated heep coin study (some spectrometers offsets changed) at 3.9 GeV KaonLT data is attached here attachment:Vijay_KaonLT_meeting.pdf . 

 ---- 

 h3. May 3, 2022 

 * Garth and Stephen's Notes attachment:mtg_22may03.odt    attachment:efficiencies_22may03.odt 
 * Ali 
 ** Fixed all the settings for high Q2 KaonLT data. Also slide on accessing spectrometer angles from CODA 
 ** attachment:KaonLT_May_3.pdf 
 * Richard 
 ** Analysis of 10.6 GeV heep data. Need to scale data to see how they compare with SIMC for smaller offsets. attachment:KaonLTMeeting5_3_22.pdf 
 * Vijay  
 ** The SIMC simulation study of the PinoLT data at Q2 = 0.38 and 0.42 GeV2 has completed (preliminary) for all settings. 
 ** An output pdf file of the 4.5 GeV, Q2 = 0.42 GeV2 and "center" setting is here attachment:Vijay_analysis_May3_2022.pdf . 

 ---- 

 h3. April 19, 2022 

 * Garth's Notes attachment:mtg_22apr19.txt 
 * Vijay  
 ** SIMC simulation updates at Q2 = 2.7, 3.6 and 4.5 GeV2. attachment:Vijay_KaonLT_Meeting_Apr19_2022.pdf   

 ---- 

 h3. April 06, 2022 

 * Garth's Notes attachment:mtg_22apr06.txt 
 * Ali  
 **    attachment:Analysis_Update_Apr_6.pdf  
 * Richard  
 ** attachment:KaonLTMeeting4_6_22.pdf 
 ** The analysis of the 10.6 GeV heep coin/singles settings are still ongoing.  
 ** There were some issues getting the replays working for the singles runs and the analysis. Things finally were sorted out yesterday (and this morning). 
 ** The SHMS heep singles are presented. Overall, they look fine. There are still some things to sort out. 
 ** The HMS heep singles are still running. 
 * Vijay  
 ** The simulation study of the 4.9 GeV KaonLT data has completed. 
 ** This energy has left, center and right settings data. All these settings look fine.  
 ** See the pdf for more info. attachment:Vijay_KaonLT_Meeting_Apri_06_2022.pdf   

 ---- 

 h3. March 23, 2022 

 * Garth's Notes attachment:mtg_22mar23.txt 
 * Richard  
 ** 10.6 GeV heep coin was analyzed. Things look pretty good so far, still some work to be done. attachment:KaonLTMeeting3_23_22.pdf 
 ** Heep singles are next up. There have been issues replaying these runs but Ali may have fixed this issue with his parameter file changes. 
 * Nathan 
 ** Did chack of the Mode 10 plots for HGC, was discrepancy between hardware and software caused by the variable phgcer_SampNSAT being 2 by default. 
 ** after changing back to 1 discrepancy resolved. Slides: attachment:Mode10Update_2022Mar23.pdf 

 --- 

 h3. March 9, 2022 

 * Garth's Notes attachment:mtg_22mar09.txt 
 * Regina March 4 meeting on pion leak through in (e,e'K+) analysis attachment:grp_22mar04.txt 
 * Richard  
 ** Vijay's scripts are working and Richard reproduced some of his previous plots. attachment:KaonLTMeeting3_9_22.pdf 
 ** The next step is to do a full analysis of the 10.6 GeV heep coin/singles data. 
 * Nathan 
 ** Update on the current state of mode 10: attachment:Mode10Update_2022Mar9.pdf  
 ** See Mark Jones' slide on changes to hcana: https://redmine.jlab.org/attachments/1424 

 --- 

 h3. February 23, 2022 

 * Garth's Notes attachment:mtg_22feb23.txt 

 --- 

 h3. January 26, 2022   

 * Garth's Notes attachment:mtg_22jan26.txt 
 * Vijay, Heep updates from the PionLT experiment. attachment:Analysis_Updates_Jan26_2022.pdf  
 ** HMS single 2.7 GeV, total 6 settings. attachment:HMS_Analysed_2p7_1.pdf attachment:HMS_Analysed_2p7_2.pdf attachment:HMS_Analysed_2p7_3.pdf attachment:HMS_Analysed_2p7_4.pdf attachment:HMS_Analysed_2p7_5.pdf attachment:HMS_Analysed_2p7_6.pdf    
 ** SHMS single 2.7 GeV,total 6 settings. attachment:SHMS_Analysed_2p7_1.pdf attachment:SHMS_Analysed_2p7_2.pdf attachment:SHMS_Analysed_2p7_3.pdf attachment:SHMS_Analysed_2p7_4.pdf attachment:SHMS_Analysed_2p7_5.pdf attachment:SHMS_Analysed_2p7_6.pdf 
 ** The 3.6 and 4.5 GeV data do not have elastic data. 
 ** Coin: 2.7, 3.6 and 4.5 GeV. attachment:Analysed_2p7.pdf attachment:Analysed_3p6.pdf attachment:Analysed_4p5.pdf    
 * Nathan, Short update on the Mode 10 
 ** Short update. May need to update hcana to change software voltage threshold. attachment:2022-01-26NathanHeinrich.pdf 
 * Jacob, EDTM Study 2 
 ** EDTM study with beam, taken during LUMI Scan 3 
 ** Early look at 3rd EDTM Study 
 ** Additional notes on increased deadtime in SHMS at High Current. attachment:PionLT_EDTM_Study2_Analysis.pdf 

 ---- 

 h3. January 12, 2022   

 * Garth's Notes attachment:mtg_22jan12.txt 
 * Junaid 
 ** PID study 
 * Vijay 
 ** Heep-singles study attachment:Analysis_Updates1_Jan12.pdf  
 * Nathan 
 ** Mode-10 data variables 

 ---- 

 h3. January 5, 2022   

 * Garth's Notes attachment:mtg_22jan05.txt 
 * Stephen 
 ** Stephen was updating GitHub but UTIL_BATCH has some permissions issues 
 * Richard 
 ** Update to KLT task schedule 
 ** Meeting between Richard and Vijay to discuss task breakdown Friday, January 7th at 11 am EST 
 ** EDTM study 2 will done sometime this weekend 
 * Vijay 
 ** Breakdown of HEEP single analysis for low Q2 data 
 *** Shapes of SIMC and data are very off 
 *** It seems to be a systematic issue on Vijay's part because the same issue is being seen in PLT data 
 *** The main issue in my input file was the target thickness, which picked up by Dave in the group meeting. I corrected the target thickness and now it looks ok. See the attached pdf file below.   
 ** https://redmine.jlab.org/attachments/1376 
 ** https://redmine.jlab.org/attachments/1375 
 * Nathan 
 ** Update on fADC mode 10 studies and the fADC deadtime studies 
 ** https://redmine.jlab.org/attachments/1372 
 * Jacob 
 ** Summary of EDTM analysis so far in the PionLT experiment 
 ** https://redmine.jlab.org/attachments/1373 

 ---- 

 h2. 2021 Meetings 

 h3. August 4, 2021 

 * Richard 
 ** Update on the HEEP singles. The remainder of energy setting ratios vs SIMC were looked at over the range of angles. They look good (within 1% and < 0.5% uncertainty). 
 ** Issues with the COIN are seen but may be a replay issue. 

 ---- 

 h3. July 21, 2021 

 * "Richard":https://redmine.jlab.org/attachments/download/1250/KaonLTMeeting7_14_21.pdf 
 ** Update on the HEEP singles. 10.6 GeV ratios vs SIMC were looked at over the range of angles. They look good (within 1%) but proper error bars are needed. 

 ---- 

 h3. May 12, 2021 

 * Richard 
 ** The currents were adjusted from 1-10 uA with little effect on the EDMTLT but proper changes to the CPULT. The replay bcm cut script developed by Sanghwa was also analyzed to see if it may be causing the issue but it had the same effects. There was also a small error in the code found where T.coin.pEDTM.tdcTimeRaw was incorrectly written as T.coin.pEDTM.tdcTime. 

 ---- 

 h3. April 28, 2021 

 * Richard 
 ** Lumi table was created for 2 sets of runs. Odd behavior seen ~30-45 uA with EDTM where is drops. Playing with the current cuts is the next step. 

 ---- 

 h3. April 14, 2021 

 * "Richard":https://redmine.jlab.org/attachments/download/1197/KaonLTMeeting_4_14_21.pdf 
 ** Updated Lumi carbon plots with uncertainties. PID and current cuts need to be played around with a bit to see if they improve things (as uncertainties don't explain any boiling-type behaviors).  
 ** Some PionLT carbon runs were looked at but there are still some issues (either cut or pathing related). 

 ---- 

 h3. March 31, 2021 

 * Richard 
 ** Luminosity studies 
 *** PionLT carbon runs 
 **** Still having issues getting decent results. I think I am just missing a definition somewhere that I need to track down 
 *** Uncertainties 
 **** Easy to do, I just need to put it into my script 
 ** Explicit PS calculations 
 *** I have gotten a rough script working for very simple runs (ie low rate, single spectrometer, calibration runs) 
 *** I am trying to expand to lumi singles runs, things get a bit more complex so I am just working out the trigger correlations 
 ** Pass 2 runs 
 *** DEF-files 
 **** I checked out the DEF-files for our combination of the nested files 
 **** There are a lot a files (~150!!!!) and may be hard to track so next week *we should make a flowchart together* 
 **** Working on my list of branches for the next replay 
 ***** Mainly just working on my list of branches for the explicit PS calculations 
 ***** *Everyone should have a list of these branches by next mini-meeting (4/7/21)* 
 * Stephen 
 ** Updated DEF-File for the "full" replay 
 *** Can't remove NData brances, Mark says they won't take up too many resources anyway 
 *** Removed individual PMT info and raw info for many detectors, reduced file size by factor ~2 
 ** Updating DEF-files/scripts for calibrations as part of PionLT prep 
 ** Will focus on DEF-file/DEF-cut file updates over next week 
 *** Will investigate HeeP replays over rest of this week. 

 ---- 

 h3. March 24, 2021 

 * KaonLT Mini Meeting focusing on Pass2 preparation 
 * "Task summary from Richard":https://redmine.jlab.org/projects/kltexp/wiki/Replay_2_Tasks 

 ---- 

 h3. March 17, 2021 

 * "Richard":https://redmine.jlab.org/attachments/download/1197/KaonLTMeeting_3_17_21.pdf 
 ** I have re-run the carbon lumi runs with the changes to the replay events (i.e. All_events_NO_EDTM). The results look pretty good (within a few percent). There are still some trends I need to investigate for certain runs and I need to do the analysis for the pionLT carbon lumi data. 
 * Mini meeting March 24th @ 4:30 EDT 

 ---- 

 h3. March 3, 2021 

 * "Richard":https://redmine.jlab.org/attachments/download/1197/KaonLTMeeting_3_3_21.pdf 
 ** I found an issue with the replay scripts used where the EDTM was not being properly registered. There is a discrepancy between two replay scripts depending on what the Decode_master is set to (i.e. ALL_events vs ALL_events_NO_EDTM).  
 ** The explicit calculation of the prescale has hit a snag because of the differences between 6 GeV and 12 GeV. I contacted Mark Jones, who explained there are significant differences in the ADC gate between the two eras (which my own findings confirmed). I have contacted Brad for further input. 

 ---- 


 h3. February 17, 2021 

 * Richard 
 ** Tried to remove some cumbersome elements to the analysis scripts without success. A few runs were added to a well behaved set of lumi runs with good consistency. 
 * Nathan 
 ** Presented an update on the Hodoscope Calibrations: https://redmine.jlab.org/attachments/download/1196/HodoscopeUpdate_2021Feb17.pdf 

 ---- 

 h3. February 3, 2021 

 * Richard 
 ** A table with the aerogel efficiency for kaons is being created. It will be expanded to other particles and detectors. 

 ---- 

 h3. January 20, 2021 

 * "Richard":https://redmine.jlab.org/attachments/download/1172/KaonLTMeeting_1_20_21.pdf 
 ** All KaonLT carbon runs have been analyzed. There are still some issues that seem to be related to the pre-scale values. I may need to investigate the EDTM and/or the COIN events further to correct for these. 

 ---- 

 h3. January 6, 2021 

 * "Richard":https://redmine.jlab.org/attachments/download/1153/KaonLTMeeting1_6_21.pdf 
 ** Analyzed most of the luminosity scans. There are a few issues I am investigating still which seem to be run related.  

 ---- 

 h2. 2020 Meetings 

 h3. December 2, 2020 
 
 * Ali and Richard 
 ** Richard included the EDTM to Lumi analysis and found some old analysis of scalers by Mark Jones that he will compare. Ali and Richard met last week and will begin combining Ali's tracking studies with Richard's lumi analysis. 

 ---- 

 h3. November 18, 2020 

 * "Carlos":https://redmine.jlab.org/attachments/download/1103/LiveTime_Study_cyero.pdf 
 ** He is investigating some of the EDTM rate dependence and sees a discrepancy between the total EDTM LT and CPU LT with a sudden dip at 20-30 kHz. More luminosity runs will need to be studied for a full conclusion. 
 * "Richard":https://redmine.jlab.org/attachments/download/1101/KaonLTMeeting11_18_20.pdf 
 ** Scaler analysis matches Carlos' code now. Overall significant improvements were made through some cut adjustments and incorrect variable definitions. 
 * Stephen 
 ** Updated the hallc_replay_lt with parameter files for all coin peaks. 

 ---- 

 h3. November 4, 2020 
 
 * "Richard":https://redmine.jlab.org/attachments/download/1092/KaonLTMeeting11_4_20.pdf 
 ** Scaler analysis is getting closer to an answer. An in-depth comparison of Carlos and Richard's analysis shows inconsistencies between the two.  

 ---- 

 h3. October 28, 2020 

 * Ali 
 ** Went over some of his tracking procedures for the report file implementation 
 * "Richard":https://redmine.jlab.org/attachments/download/1091/KaonLTMeeting10_28_20.pdf 
 ** Luminosity analysis continues. There are still some issues but things have improved a bit. 
 * "Stephen":https://redmine.jlab.org/attachments/download/1090/KaonLT_Meeting_28_10_20.pdf 
 ** Discussion of CT peak analysis, progress and outline of what's left to do with this 
 * Vijay 
 ** Dry run for Mark's hallc meeting tomorrow (10/29) 

 ---- 

 h3. October 14, 2020 

 * Ali 
 ** Working through the REPORT_FILES template to get everything working and outputting proper information. 
 * "Richard":https://redmine.jlab.org/attachments/download/1085/KaonLTMeeting_10_14_20.pdf 
 ** Updates on lumi scans. There are many issues from Scalers to Tracking that need to be resolved. 
 * Stephen 
 ** He has included a demo for using the python analysis script in the README. 

 ---- 

 h3. September 23, 2020 

 * "Richard":https://redmine.jlab.org/attachments/download/1078/KaonLTMeeting_9_23_20.pdf 
 ** A quick lumi analysis after the Pass1 runs. The results were compared to previously analyzed runs (5154-5158). There was some change but nothing huge. Some other runs with the same energy settings were analyzed as well.  
 * Stephen 
 ** Showed some slides from u-channel meeting of omega and proton MM    plots. Then discussed some ideas for dealing with the hole in the HGCer. 
 * Petr 
 ** Data fitting algorithm based on E.H. Bellamy paper. Extracts PMT spectra from ROOT file and makes a 1D projection histogram. 

 ---- 

 h3. September 9, 2020 

 * Richard 
 ** Still working on run list, GUI checks 
 * "Stephen":https://redmine.jlab.org/issues/567 
 ** New slim replay scripts and full replay scripts updated 
 * Ali 
 ** Him and Stephen found an issue (already aware of) with zeros in the parameter files using Mark's hcana changes 
 ** Need to be vigilant of crashed replays because this is probably the issue 
 * Vijay 
 ** Nearly finished with HGcer calibrations 

 ---- 

 h3. August 26, 2020 

 *Links to talks->* "Ali":https://redmine.jlab.org/issues/483 "Vijay":https://redmine.jlab.org/issues/479 

 * Richard 
 ** New tasks were created based on the previous discussions on objectives going forward. In the process of checking run sheets against the run list and camera/EPICS. 
 * Stephen 
 ** Updated on RF times. There were a few issues with some runs, but overall most are finished. 
 * Ali 
 ** Hodoscope calibrations look similar to before. There is still the issue of the double bands in the time walk. This is ok because of the wide windows we still have. (See "hodoscope calibration section":https://redmine.jlab.org/projects/kltexp/wiki/Analysis_Tasks#Calibrations-Mid-to-Late-August-2020) 
 * Vijay 
 ** Updated the SHMS HGCher analysis. The calibrations are coming along, but a little more work needs to be finished up. 

 ---- 

 h3. July 29, 2020 

 *Links to talks->* "Richard":https://redmine.jlab.org/attachments/download/1029/klt_7_29_20.html "Stephen":https://redmine.jlab.org/documents/106 

 * Stephen 
 ** Updates on scripts to set pedestal defaults and AdcTdc offsets after hcana change 
 * Richard 
 ** Updated plan and schedule for the analysis going forward 

 ---- 

 h3. July 8, 2020 

 *Links to talks->* "Vijay":https://redmine.jlab.org/attachments/download/1023/SHMS%20Proton%20Absorption.ods "Ali":https://redmine.jlab.org/attachments/download/1022/KaonLT_Meeting_Jul_8.pdf 

 ---- 

 h3. June 24, 2020 

 *Links to talks->* "Vijay":https://redmine.jlab.org/attachments/download/1014/Vijay_24_June_20.pdf "Richard":https://redmine.jlab.org/attachments/download/1013/KaonLTMeeting_6_24_20.pdf 

 ---- 

 h3. June 9, 2020 

 *Links to talks->* "Vijay":https://redmine.jlab.org/attachments/download/1006/Vijay_9_June_20.pdf "Ali":https://redmine.jlab.org/attachments/download/1007/KaonLT_Meeting_Jun_9.pdf "Richard":https://redmine.jlab.org/attachments/download/1005/KaonLTMeeting_6_9_20.pdf 

 ---- 

 h3. May 27, 2020 

 *Links to talks->* "Vijay":https://redmine.jlab.org/attachments/download/1002/Vijay_26_May_20.pdf "Ali":https://redmine.jlab.org/attachments/download/1003/KaonLT_Meeting_May_27%20%5BAutosaved%5D.pdf 

 ---- 

 h3. May 13, 2020 

 *Links to talks->* "Stephen":https://redmine.jlab.org/attachments/download/993/KaonLT_Meeting_13_05_20.pdf - Python analysis scripts "Richard":https://redmine.jlab.org/attachments/download/998/KaonLTMeeting_5_13_20.pdf "Ali":https://redmine.jlab.org/attachments/download/997/KaonLT_Meeting_May_13.pdf "Vijay":https://redmine.jlab.org/attachments/download/999/Vijay_13_May_20.pdf 

 ---- 

 h3. April 29, 2020 

 *Links to talks->* "Richard":https://redmine.jlab.org/attachments/download/984/KaonLTMeeting_4_29_20.pdf "Stephen":https://redmine.jlab.org/attachments/download/985/KaonLT_Meeting_29_04_20.pdf - Python code planning "Mark's hcana tracking update":https://redmine.jlab.org/attachments/download/982/Updates-to-HCANA-tracking.pdf 

 ---- 

 h3. April 15, 2020 

 *Links to talks->* "Richard":https://redmine.jlab.org/attachments/download/977/KaonLTMeeting_4_15_20.pdf 

 ---- 

 h3. April 1, 2020 

 *Links to talks->* "Stephen":https://redmine.jlab.org/attachments/download/949/KaonLT_Meeting_01_04_20.pdf - RF timing "Richard":https://redmine.jlab.org/attachments/download/948/KaonLTMeeting_4_1_20.pdf 

 ---- 

 h3. March 18, 2020 

 *Links to talks->* "Richard's EIC Temple Draft 2":https://redmine.jlab.org/attachments/download/942/DRAFT2_mesonSF_EIC_Temple.pdf 

 ---- 

 h3. March 11, 2020 

 *Links to talks->* "Ali":https://redmine.jlab.org/attachments/download/940/KaonLT_Meeting_Mar_11.pdf "Richard's EIC Temple Draft 1":https://redmine.jlab.org/attachments/download/941/DRAFT1_mesonSF_EIC_Temple.pdf 

 ---- 

 h3. Feb 26, 2020 

 * Peter's changes to hcana should be coming soon. Richard will check with Mark when he is down there for an update. 
 * Need to incorporate the latest "BCM updates":https://hallcweb.jlab.org/doc-public/ShowDocument?docid=1034 from Dave Mack. 
 * Richard should look at the luminosity scalar analysis before and after the BCM changes once he gets his script working again. 
 **    He is checking with Mark about this while he is down in jlab.  

 ---- 

 h3. Feb 10, 2020  

 ---- 

 h3. Jan 22, 2020 

 *Links to talks->* "Richard's Hall C Winter Talk Draft 1":https://redmine.jlab.org/attachments/download/937/KaonLTHallC_WinterTalk2020_Draft1.pdf 

 ---- 

 h3. Jan 15,2020 

 *Links to talks->* "Richard's Hall C Winter Talk Outline":https://redmine.jlab.org/attachments/download/936/KaonLTHallC_WinterTalk2020.pdf 

 ---- 

 h3. Jan 8,2020 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/76 

 * Richard 
 ** PID efficiency script is coming together 
 *** Tighter cuts and cuts on timing, beta, etc. will be needed for cleaner electron cuts 
 *** Hcal cut should be changed to cut around the mean normalized value instead of cutting above or below it only. This will assure electrons aren't lost. 

 ---- 

 h2. 2019 Meetings 

 ---- 

 h3. Dec 17,2019 

 Reminder: Mark Jones' commissioning analysis meetings are bi-weekly and announced by Mark 

 * Stephen 
 ** Number of passes finished for HMS calorimeter 
 *** Still wiggles in all passes 
 ** Finished his report on calibrations, "DC and Calorimeters":https://redmine.jlab.org/documents/74 
 ** Put together a list of people's jobs from Regina over the next couple of months, "Regina Analysis Tasks":https://redmine.jlab.org/documents/75 
 *** CUA will be putting together a similar list 

 * Richard 
 ** Fixed one error issue with SHMS calorimeter calibration code (discussed Dec 10th meeting), but still have one error 
 ** PID efficiency script plots are complete, but still some questions he needs to work out before a final version 
 *** Eventually need to include tracking into PID efficiencies 
 ** Pete brought up the question for a number of multi-hits we are getting 
 *** (Dead Time)+(Trigger efficiency)-EDTM 

 ---- 

 h3. Dec 10,2019 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/73 

 * Garth and Stephen were in the most recent SIDIS meeting 
 ** Wit had lots of hcana changes but they seem to be critical to our work  
 *** Changes such as new detector classes 
 ** Hopefully Wit will present these in the next commissioning meeting 

 * Stephen improved wiggles but still prevalent 

 * Richard finished HGCer calibrations, he concluded that the online parameters are sufficient 
 ** He also has been looking at SHMS calorimeter calibrations but has had issues running the scripts. Most likely a pathing issue. 
 ** The "tasks":https://redmine.jlab.org/projects/kltexp/wiki/Analysis_Tasks page has been updated with a section for efficiencies.  
 *** A PID efficiency script is currently being developed and hopefully should have some plots for our next meeting. 

 ---- 

 h3. Nov 26,2019 

 *Links to talks->* "Richard":https://redmine.jlab.org/issues/480 

 * Stephen 
 ** DC SHMS, small-angle has poor performance 
 *** Tighter cuts or PID (currently only looks for electrons) 
 ** Vardan said calorimeter wiggles are an ongoing issue with no current solution 

 * Richard 
 ** The calibrations are nearing completion with only the HG Cherenkov and SHMS calorimeter yet to be finished. There are also smaller issues with some DC and HMS calorimeter runs. 
 *** Richard will pick up looking at the HGCer calibrations from Vijay's scripts.  
 *** Richard will also be looking at the SHMS calorimeter and with the suggestion of getting some negatively polarized SHMS runs from Hem. 
 ** The HMS Cherenkov is finished with the conclusion that the calibrations from the online runs are good 
 ** We can begin looking forward towards efficiencies.  
 *** Garth suggested that looking at PID may be an advantageous first steps. Precise definitions of these PIDs will need to be logged. 
 *** Clear notes on efficiencies steps will be needed. They will need to be redefined multiple times in multiple ways.  

 * Hem 
 ** Has been looking at the calorimeter as well.  
 ** He has seen the second band in the spectra for some runs, we suggested further looks into HMS hodoscope and subsequently reference times. A recalibration will need to be performed.   

 ---- 

 h3. Nov 19,2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/issues/500 "Richard":https://redmine.jlab.org/issues/500 

 * Richard 
 ** Finished up TDC cut paper going through the procedure (found "here":https://redmine.jlab.org/issues/500) 
 ** Needed to add Carlos pre-trigger designs to an appendix for the paper which clarifies the introduction section 

 * Stephen 
 ** Issues with DC calibration for only a few runs during each run period 
 ** "wiggles" in HMS calorimeter calibrations for some runs 
 *** few runs where HMS cal totally failed (which was dependent on initial parameters) 
 *** Spring looked worse than the rest (surprisingly), but this may be because online calibration wasn't great 

 ---- 

 h3. Nov 5,2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/issues/500 

 * Timing update (Stephen) 
 ** Reference times between Richard and Stephen are very similar and the GitHub is updated accordingly 
 ** Richard is going to add the trigger time cuts once the fall runs are finished (by end of week) 
 ** First pass of detector reference time parameter files finished 

 * Peter's hcana code (Richard and Stephen) 
 ** Peter's code does not remove ghost bands (already known), but now Stephen is seeing a strang horizontal band for just SHMS hodoscope 2D plots 
 *** Richard has emailed Peter about the issue with no response yet (follow-up soon) 
 ** This means we still need the reference time amplitude cut with Peter's code but using just the amplitude cut with Mark's latest version of hcana should be sufficient for calibrations. 

 * Calibrations were decided to be done once per shift 

 * Next meeting schedule... 
 ** Small meetings Nov 12 and Nov 19 (at 3pm EST) so Richard and Stephen can iron out the timing and calibrations 
 ** November 26 (at 3pm EST) will be a larger meeting.  
 *** This is the rough deadline we made for finishing up calibrations.  
 *** This should be sufficient for all detectors (although hodoscope and HG which may persist). The ultimate goal is for all calibrations to be finished up before Winter break!! 

 ---- 

 h3. Oct 22,2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/issues/500 "Richard":https://redmine.jlab.org/issues/500 

 * SHMS Reference Time Amplitude Cuts(Stephen) 
 ** Extra peaks in HG from cross-talk and "ghost-bands" due to ADC screw up 
 *** cuts applied on SHMS ADC ref time amplitude (greater than zero) done before any further event selection and histogram filling 
 ** This is a temporary fix to "ghost-bands" and cross talk 
 *** Although for HG there is still a small peak. The hodoscope is fine. 
 ** The time window is easier to set now. 
 ** DC has been fine (since it uses TDC) 
 *** DC calibrations should be good to complete, other detectors need more fixes (i.e. the remaining "ghost-band") 
 ** Hodoscope still bad but this might be because of that "ghost-band", needs to be further investigated 

 * Peter's hcana fixes for reference time (Richard) 
 ** Peter implemented changes to hcana 
 ** Two major issues... 
 ## Amplitude goes to null and integral is low or negative which causes a rate-dependent efficiency.  
 *** Peter has used the average pulse value acquired over a large sample of runs for each detector.  
 ## Timing offset between EL-REAL and 3/4 is 130 ns which causes junk tracks at high SHMS rates (~6% of the time) 
 *** Peter created a routine to find the best offset by means of matching TDC and ADC 
 ** Richard has been checking for bugs in the code.  
 *** It seems to be ready to implement, Richard emailed Peter for confirmation on the implementation in our group directory 

 * Calibration of HMS detectors when the timing is agreed upon by Stephen and Richard. 
 ** Besides hodoscope which still has an issue with the script 
 ** Stephen and Richard are having a meeting late next week to determine time window values 
 *** These will be relatively final, but on the wider side if adjustments need to be made in the future 

 ----- 

 h3. Sep 24,2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/issues/500 "Richard":https://redmine.jlab.org/issues/500 

 ----- 

 h3. Sep 10,2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/documents/69 "Ali":https://redmine.jlab.org/documents/72 "Richard":https://redmine.jlab.org/issues/500 

 ----- 

 h3. August 26, 2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/documents/69 "Ali":https://redmine.jlab.org/documents/70 "Vijay":https://redmine.jlab.org/documents/71 "Richard":https://redmine.jlab.org/issues/500 

 ---- 

 h3. June 21, 2019 

 *Links to talks->* "Hallc Talk Outline":https://redmine.jlab.org/documents/60  

 ---- 

 h3. June 14, 2019 

 *Links to talks->* "Hallc Talk Outline":https://redmine.jlab.org/documents/60 "Mireille":https://redmine.jlab.org/documents/64 "Update-Mireille":https://redmine.jlab.org/documents/65 "Vijay":https://redmine.jlab.org/attachments/download/814/Analy_update3.pdf 


 ---- 

 h3. June 4, 2019 

 *Links to talks->* "Stephen":https://redmine.jlab.org/documents/56 "Nathan":https://redmine.jlab.org/documents/57 "Vijay":https://redmine.jlab.org/documents/58 "Richard":https://redmine.jlab.org/documents/59 "Hallc Talk Outline":https://redmine.jlab.org/documents/60 "Mireille":https://redmine.jlab.org/documents/61 


 ---- 

 h3. May 21, 2019 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/52 "Mireille":https://redmine.jlab.org/documents/? "Stephen":https://redmine.jlab.org/documents/53 "Nathan":https://redmine.jlab.org/documents/54 "Vijay":https://redmine.jlab.org/documents/55 

 ---- 

 h3. May 14, 2019 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/48 "Carlos' slides":https://redmine.jlab.org/documents/49 "Mireille":https://redmine.jlab.org/documents/50 "Stephen":https://redmine.jlab.org/documents/51 

 ---- 

 h3. April 30, 2019 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/45 "Richard's slides in wiki":https://redmine.jlab.org/projects/kltexp/wiki/Analysis_How-To "Brad Computing":https://redmine.jlab.org/documents/46 "Stephen":https://redmine.jlab.org/documents/47 

 ---- 

 h3. February 4, 2019 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/43 "Stephen":https://redmine.jlab.org/documents/44 

 ---- 

 h2. 2018 Meetings 

 ---- 

 h3. August 21, 2018 

 *Ongoing Tasks->* #381 #382 #383 #404 #405 #406 

 ---- 

 h3. August 8, 2018 

 *Ongoing Tasks->* #381 #382 #383 #404 #405 #406 

 ---- 

 h3. August 3, 2018 

 *Ongoing Tasks->* #381 #382 #383 
 *Links to talks->* 

 ---- 

 h3. July 11, 2018 

 *Ongoing Tasks->* #381 #382 #383 #384 

 *Action items:* 

 * Update run plan, e.g. add prescales (Garth) 
 * Check on details of: magnet cycling procedure, minimum angles, MCC fringe field steering, beam energy measurements, optics checks (Tanja) 
 * Identify additional run coordinators (Tanja) 
 * Encourage collaborators to take shifts, e.g. FIU (Pete) 
 * Draft email about shifts etc. to collaboration (Tanja) 
 * Develop diagnostic scripts and kaonLT specific (Richard, Ryan) 
 * Continue documentation development, e.g. checklist, shift taker instructions, HowTo or link to generic Hall C HowTo, task list, web page (Richard, Garth) 

 *Discussions:* 

 * SHMS Aerogel Cherenkov: tray has been changed to SP-11 on 10 July, nominal time for change is 8 hours, one tray can stay in the hall to minimize overhead 
 * HGC: waiting for re-installation 
 * Trigger checkout: main items to keep in mind: fADC thresholds, timing HMS/SHMS - leading arm, possible rate limitations 
 * Shifts 
 * Diagnostics scripts and documentation 
 * Personnel schedules 

 ---- 

 h3. June 28, 2018 

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/34 

 *Action Items:* 

 * Develop and post shift schedule. Send out request. Find out who is KAONLT liaison (Tanja) 
 * Contact potential run coordinators. (Tanja) 
 * Aerogel tray exchange for start of experiment and coordination of future exchanges (Vladimir) 
 * Update runplan and prepare documents needed for shift crew, e.g. run sheets (Garth) 
 * Set up a web page to collect and link from Redmine - How-To documents, initial calibrations, etc. (Garth, Richard) 
 * Adapt luminosity scan script to calculate kaon yields by run number for use during experiment by shift crew - has to be simple to use. (Richard) 
 * Develop trigger/detector checkout plan (Tanja, initial draft) 
 ** Are all experts available at run start, experts needed: coincidence timing , Harp Scans, BCM calibrations, ... 
 * Find out from JLab what the anticipated settings are for limit switches for both spectrometers and for separation angle in between them. Also find out what magnet cycling procedure should be used for L/T quality separation data (Tanja) 
 * Find out if all aerogel trays needed for experiment can be stored inside Hall C to avoid having to open truck ramp (Vladimir) 

 *Analysis Updates:* 

 * Ryan  
 ** resolving differences between his analysis and Richard's 

 * Richard 
 ** Discussed with Mark Jones and is cross checking his January luminosity scan analysis against Mark's. Realized that BCM calibrations (gains, offsets etc) are different - currently working on fixing that, presents current results - low curent point is ~10% below unity, discussion that this seems to indicate that BCM offsets are the issue as low current points would be most sensitive, is planning to also double-check the BCM calibration procedure - had discussion with Dave Mack at JLab 
 ** Presents a script developed for luminosity scans that determines yields for a run number input on command line. Discussion this could be good to adapt for kaon counting during experiment. 

 ---- 

 h3. May 31, 2018 

 *Links to talks->* "Samip":https://redmine.jlab.org/documents/30 

 ---- 

 h3. May 17, 2018  

 *Links to talks->* "Richard":https://redmine.jlab.org/documents/29, "Ryan":https://hallcweb.jlab.org/DocDB/0009/000968/001/Presentation.pdf, "Samip":https://redmine.jlab.org/documents/30 

 * Luminosity scans (Richard, Ryan) 
 ** Carbon scan from April seems consistent with no boiling, LH3 and LD2 show deviation from unity, should check effect of EDTM events on efficiencies 
 ** Carbon scans from January (low rates) have large deviations from unity, in particular behavior at higher currents (>40uA), e.g. cpu dead time, that is not yet understood 
 ** discussion about pre-scale factors 

 * Discussion about Richard's script extracting and printing information from report files and trees including histograms 
 ** Add option for looking at yields with cuts on cointime, target window, and missing mass 
 ** Add option to look at real peak vs. background level 

 * H(e,e'p) kinematics investigation for fall run (Samip) 
 ** Optimal range is between electron angles of 17-20 degrees, time required ranges from ~1 hour at 17 degrees to ~3 hour at 20 degrees, other angles may be of interest for disentangling charge and current dependent effects 

 * *Action Items* 
 ** Luminosity scans at low rates from January (Richard) 
 ** Luminosity scans from April with corrected efficiencies, e.d. DT (Ryan) 
 ** H(e,e'p) and H(e,e') kinematics for upcoming Kaon LT run (Samip) 
 ** Kaon LT detailed run plan (Garth, Pete, Tanja) 
 ** Preparations, e.g. report files, diagnostic histograms and real time monitoring, shift worker instructions, optimization of trees (all) 

 ---- 

 h3. May 10, 2018 

 *Links to talks->* "Hem":https://redmine.jlab.org/documents/24, "Richard":https://redmine.jlab.org/documents/23, "Ryan, ":https://redmine.jlab.org/documents/20 "Samip":https://redmine.jlab.org/documents/26 

 * Luminosity Scans January 2018 (Richard) 
 ** Tracking-based analysis for Runs 1415-1423, .. for now focus on HMS, typical HMS S1X rates are < 
 ** Cuts applied: PID for electrons (Cer, Cal), acceptance (delta) 
 ** Efficiency calculation was investigated in detail and applied using scalers from trees 
 ** Results for carbon (normalized to 30uA) show large variation - need to understand 
 ** *Next steps:* 
 *** Double check calculation of computer lifetime, look for differences in TRIG and PRETRIG scalers 
 *** Check stability of pre-scale (Ryan saw changes within runs) 
 *** Check if detector calibrations for runs are implemented correctly 
 *** Analyze with Ryan's cuts 

 * Luminosity Scans April 2018 (Ryan) 
 ** Confirms Richard's earlier finding that scaler-based analysis is not reliable and that it's best to analyze event-by-event 
 ** Tracking-based analysis for runs 2073-2095 
 ** Cuts applied: PID for electrons (TOF, Cal, Chi2), acceptance (delta and angles), BCM4a 
 ** Included most recent detector calibrations in replay of these runs 
 ** Efficiencies from report file mainly, may need to update tracking efficiency implementation 
 ** Found that calculations of computer deadtime requires different set of scalers for runs from January to those from April 2018 
 ** Results show carbon yields consistent with unity within error (0.5% error bars), and boiling for hydrogen and deuterium 
 ** *Next steps:* 
 *** Check tracking efficiency 

 * *Global comments on Luminosity Scans (so far)* 
 ** April 2018 scans show hydrogen and deuterium target boiling on level of 3-4%, Carbon is consistent with no boiling, as expected - based on description, new targets are expected to have less boiling than those used previously, e.g. in FPI2. Double check this. 
 ** Efficiencies in report files need to be updated, e.g. CPU deadtime (Richard) 
 ** Check if BCM4a is the one closest to the target - get a map of the hall (Richard, Ryan) 
 ** Uncertainties on current 
 ** Detector calibrations 
 ** HMS vs. SHMS 

 * eK+ coincidence time from SIDIS Spring 2018 run (Samip) 
 ** Worked with Mark Jones on coincidence time calculation - now available in hcana (see Samip's 10 May 2018 email for more details) 
 ** runs used in this analysis: 3423 with Q2=3.1 GeV2 and W=2.78 GeV 
 ** Coincidence time spectrum shows peak separations >2 ns corresponding to different accelerator frequency 
 ** Missing mass spectra: Kaon peak clearly visible, also see pions even after cuts 
 ** Found that pions appear localized in phase space - seems to be due to end point in kinematic phase space 
 ** *Next steps:* 
 *** Look for a run with kaon momentum closer to 3 GeV/c (aerogel threshold is ~3 GeV/c for kaons for n=1.015) 

 * Yield normalized ratios from SIDIS (Hem) 
 ** Yield ratios of data to MC match well in shape, but normalization seems not quite right 
 ** *Next steps:* 
 *** Check if efficiencies, currents, run time etc. are taken into account correctly 

 ---- 

 h3. April 26, 2018 

 *Links to talks->* "Richard,":https://redmine.jlab.org/documents/18 "Ryan, ":https://redmine.jlab.org/documents/20 "Samip":https://redmine.jlab.org/documents/25 

 * Discussion luminosity scan analysis (earlier scan) - Richard 
 ** Shown is a scaler-based analysis using scalers from report file and from ROOT tree - the values are different, why that is has to be checked 
 ** Notes on efficiencies in report file: 
 *** DT (EDTM) is not electronic deadtime. Instead use for now 50ns* trigger rate for HMS and 100ns* trigger rate for SHMS 
 *** Cherenkov efficiency seems low (check if that is due to PID, thresholds, pi/e, etc) 
 *** Tracking efficiency needs to be checked 
 *** 3/4 trigger efficiency seems ok 
 * Discussion luminosity scan analysis (later scan) - Ryan 
 ** Fewer carbon runs than earlier scan 
 ** Shown is a tracking-based analysis using data from the ROOT trees 
 ** First results for carbon, LH3, and LD2, but efficiencies need to be checked 
 * SIDIS analysis - Hem 
 ** Shown is a data/MC comparison for different kinematic settings. Agreement in shape is very good. Normalization has yet to be applied. 
 ** Discussion about agreement in z and apparent offset - relation of this offset with spectrometer offset 
 * Analysis of exclusive events in SIDIS - Samip 
 ** Shown is a look at the first kaon events in exclusive limit of SIDIS data 
 ** Ongoing work on implementing coincidence time into tree 
 * Discussion of run plan, and in particular possible low Q2/low energy 
 ** Assume that kaon low Q2=0.4 already included - have kinematics with specified beam energies, if needed 
 ** For pion form factor and LT, have two possible settings that can use the prescribed beam energies: Q2=1.45 and 0.35 GeV2 - prioritize towards the higher Q2 as not as sensitive to offsets 
 ** Garth will put this together in a draft email 

 ---- 

 h3. March 29, 2018 

 *Links to talks->* "Richard, ":https://redmine.jlab.org/documents/16 "Samip":https://redmine.jlab.org/documents/27 

 * Details of adding points to fall run plan - will be ready sooner, likely in week of 4/1 (Garth, Samip, Tanja) 
 * Luminosity scans (Ryan, Andres, Richard, Pete) 
 * Macros, SIDIS analysis (Hem, Anashe) 
 * Elastic data (TBD) 

 ---- 

 h3. February 22, 2018 

 *Links to talks->* "Andres":https://redmine.jlab.org/documents/19 

 *    Discussion of Andres' slides on elastic W calculation in hcana - initial issue with negative W values seems to be due to outdated information in standard.kinematics file. Additional investigation ongoing. 
 * Brief description of Samip's work - working on coincidence timing decoding with JLab staff and re-optimization of kinematics 
 * Discussion of Richard's slides on run plan kinematics and preparations - 
 ** everyone agrees that in general want to have as many L/T cross section separations as possible in a robust run plan balancing statistics vs. systematics 
 ** before actual run may want to do a 3D optimization in all kinematic variables to get the best possible delta(epsilon) for the L/T separations 
 ** for each setting will need 1-2 shifts for calibrations - need to negotiate with other experiments in run group 
 ** kinematics plots W-Q2 show that still need to do some optimization 
 ** run sheets: include raster size, update coincidence count labeling, and carbon target labeling 
 * Discussion of next meeting. Garth at JLab 8-16 March - might have meeting during that time. 

 * *Action Items:* 
 ** Post slides/notes from this meeting and procedure to set up hcana in Redmine (Tanja) 
 ** Update run sheet and kinematic optimization for W/Q2 (Garth) 
 ** Circulate elastic calculations for kinematic settings (Samip) 
 ** Continue investigation of elastic data and run plan kinematics (Andres/Richard) 
 ** Test software installation instructions (Pete) 
 ** Draft email to JLab management about need for small angle beam pipe - need to have HMS minimum angle of less than 13 degrees for kaon L/T separations (Tanja)  

 ----